Wendy was awarded a volleyball scholarship to the University of Michigan, so on graduation her parents gave her the $14,000 they had saved for her college tuition. She opted to invest some money in a privately held company that pays 10% per year and evenly split the remaining money between a market account yielding 2% and a high rick stock that yeilded 40%. At the end of the first year she had $16,610 total. How much did she invest in each of the three?
Can I please get some help on this because I have a test tomarrow and I don't understand these problems.
I've tried : .4x+(14000-2x)+.02(14000-(x+y)=16610 but I stopped when I got x+y=4147222.22 and I'm really confused now. Can you please show me how to solve this problem?

Answer by scott8148(6628)

y + 2x = 14000 ___ y = 14000 - 2x

(10%)y + (2%)x + (40%)x = 16610 - 14000

solve using substitution